# Trace: Contrastive learning for multi-trial time-series data in neuroscience

### v1

@@ 6 Conclusion, limitations, and future work
We presented Trace, a new framework that combines contrastive learning with neighbor embeddings to directly generate interpretable 2D visualizations of large-scale neural time-series data. Using the inherent structure of multi-trial recordings common in neuroscience experiments, Trace is able to separate subtle differences of simulated neural response types that competing methods fail to distinguish. When applied to a diverse neural dataset of two-photon recordings, Trace captured both continuous variations in neural properties and discrete cell-type structures. Additionally, Trace identified clusters with fine differences in functional responses and highlighted recording artifacts in the learned 2D representation (Sec. 5.3).
@@
Conceptually, Trace has two advantages over existing methods. First, methods like $t$-SNE and UMAP based on neighborhood relationships in the high-dimensional space are sensitive to uninformative, noisy baseline activity and may not be able to detect short class-informative response period (Sec. 5.1). Contrastive frameworks such as CEED rely on general-purpose data transformations and may mask subtle response differences between cell types. Second, Trace directly produces a 2D visualization while other contrastive methods require a separate reduction step. As a result, the embeddings produced by Trace were the most biologically interpretable as revealed by their ability to reflect key response characteristics (Sec. 5.2).
@@
A limitation of our work is the need to train an embedding model, increasing computational time compared to $t$-SNE or UMAP. However, this comes with the benefit of improved embedding quality and the ability to map new data points. Another limitation of our work is the need for repeated trials. However, this data collection approach data is standard for many neuronal time-series datasets.
In future work, it will be interesting to apply Trace to a much wider range of such data including spike train data following more Poisson-like statistics [34]. Here, the data augmentation approach of Trace to average subsets of trials could be particularly beneficial, as standard data augmentations such as the addition of Gaussian noise do not respect the statistical structure of the spike train.
@@ Appendix A Experimental details
A.1 Neural Data classes
A typical coarse classification of retinal ganglion cells is by their response to On- and Off-sets in the light stimulus and can typically be obtained with high confidence from simple statistics of the mean trial response of the moving bar stimulus [5]. In particular, suppressed-by-contrast responses were identified as the ones with negative area under the curve, calculated from their mean response across all directions of the moving bar. ON-OFF responses were identified using the characteristic double-peak (2nd component in PCA captured this double-peak). Finally, ON and OFF responses were classified based on their response direction (captured in 1st component of PCA).
@@ A.2 ON-OFF index
Then, metric was clipped within the range $[-1,1]$ to obtain a bounded metric.
@@ A.4 Outlier detection
A.5 Implementation details
Trace uses a lightweight multi-layer perceptron (MLP) that can run efficiently on a GPU. To ensure a fair comparison, we matched CEED’s architecture [40] and used an MLP consisting only of four layers with sizes [768, 512, 256, 128] and ReLU activations between them. We added a 1024-dimensional projection head. Due to TS2Vec’s specialized augmentations and hierarchical contrastive scheme, harmonizing its architecture was more difficult and we stuck to its original, dialated CNN architecture.
For the neural dataset, we used optimal hyperparameters found with a grid search with batch sizes ranging from 1024 to 3200 and learning rates from 0.1 to 0.2 for both Trace and CEED. The best hyperparameter setting was chosen based on the final loss. For Trace the best batch size was 1280, while for CEED it was 1024. A learning rate of 0.1 was optimal for both. For Trace $+$ CEED we used hyperparameters as in the unmodified Trace version.
@@ A.5 Implementation details
Unless otherwise specified, we trained all embeddings for 1000 epochs for the neural dataset.
@@
A.6 CEED augmentations
Here we give details on the general-purpose time-series augmentaions that we used when running CEED and Trace$+$CEED. We apply these augmentation to the mean of all trials,
@@ A.6 CEED augmentations
$y_{i}=r\cdot x_{i}\,\text{ and }\,y_{i^{\prime}}=r^{\prime}\cdot x_{i}\quad\text{ with }r,r^{\prime}\sim\mathcal{U}(0.7,1.3),$ (10)
@@
$y_{i}[t]=x_{i}[t-s]\,\text{ and }\,y_{i^{\prime}}[t]=x_{i}[t-s^{\prime}]\quad\text{ for }s,s^{\prime}\sim\mathcal{U}\big{(}\{\pm 1,\pm 2,\pm 3\}\big{)}.$ (11)
To efficiently handle the computationally expensive correlated noise generation, we pre-computed 10,000 noise samples from the covariance matrix of the neural data using a multivariate normal distribution and added them to the original time series, mimicking the augmentation
$y_{i}=x_{i}+\varepsilon\,\text{ and }\,y_{i^{\prime}}=x_{i}+\varepsilon^{\prime}\quad\text{ for }\varepsilon,\varepsilon^{\prime}\sim\mathcal{N}\big{(}\mathbf{0},\text{Cov}(\{x_{1},\dots,x_{n}\}\big{)}.$ (12)
We applied each transformation independently, with probabilities of 0.7 for amplitude jittering, 0.6 for temporal jittering, and 0.5 for correlated noise injection. Hyperparameters of the transformations were adapted to our dataset.
A.7 Supplementary tables and figures
Figure S1: Trace is able to separate classes even for high noise levels and outperforms other methods. (b) Trace embedding for increasing amounts of noise during the neural baseline activity. Discriminability (D) values indicated at the top of each embedding. Standard deviation (SD) of noise is indicated at the top of the figure. (b) Same as (a), but for CEED, (c) TS2Vec, (d) $t$-SNE, and (e) UMAP.
Table S1: Mean discriminability for simulated responses across three seeds for different noise levels (standard deviation of baseline response).
@@ A.7 Supplementary tables and figures
Figure S2: Trace embeddings with varying numbers for trial-averaging. (a) From left to right: Embeddings for the synthetic dataset with increasing numbers of trials $k$ from 1 to 7 used for the non-overlapping subsets. Discriminability (D) between class 1 and class 2 neurons was computed as in Fig. 2. Bottom: Loss over epochs. (b) From left to right: Embeddings for the neural dataset with increasing numbers of trials per subset mean $k$ from 1 to 7 for the ”chirp” stimulus and from 1 to 5 for the moving bars stimulus. Neurons are colored by the four broad groups OFF (blue), ON-OFF (turquoise), ON (green), Suppressed-by-contrast (black). Bottom: Loss over epochs.
Table S2: Reducing number of trials ($k$) used for the non-overlapping subsets using the neural dataset. The maximum number of trials was 7 for the ”chirp” stimulus ($k_{\text{chirp}}$) and 5 for the moving bar stimulus ($k_{\text{moving bar}}$).
@@
Figure S3: Trace-variants visually exhibit a good balance between resolving finer cluster structure and retaining large-scale structure. (a – c) Two-dimensional embeddings of (a) Trace, (b) Trace+CEED, (c) Trace+TS2Vec. Color-coded according to their functional group.
@@
Figure S5: Identified artifacts across all methods. (a) Trace embedding with identified artifact island in red. (b) Same as in (a) but for Trace $+$ CEED, (c) Trace $+$ TS2Vec, (d) CEED, (e) TS2Vec, (f $t$-SNE, and (g) UMAP.
@@
Table S3: Evaluating dimensionality reduction options for visualizing CEED’s 5-dimensional embeddings in 2D space. For the different normalizations, see the caption of Fig. S6.
# Model $r_{\text{OOi}}$ $r_{\text{RTi}}$ $r_{\text{Depth}}$
1 CEED (non-norm PCA) 0.58 0.14 0.54
2 CEED (non-norm UMAP) 0.36 0.19 0.54
3 CEED (unit-norm PCA) 0.55 0.11 0.55
4 CEED (unit-norm UMAP) 0.38 0.16 0.31
5 CEED (z-norm PCA) 0.59 0.12 0.53
6 CEED (z-norm UMAP) 0.42 0.09 0.39
Table S4: Training time for all methods. Data augmentations for CEED and TRACE $+$ CEED were pre-computed and sampled from during training to improve training time (see Sec. A.6). Please note that no data augmentations are necessary for Trace. Each experiment was run on a NVIDIA A40 GPU 48 GB. For implementation details, see Sec. A.5.
@@
Trace $+$ CEED 1000 25:40 $\pm$ 5.1s
Trace $+$ TS2Vec 1000 962:43 $\pm$ 2668.8s
CEED $\xrightarrow{}$ PCA 1000 25:17 $\pm$ 19.7s
TS2Vec $\xrightarrow{}$ PCA $<$1 1:23 $\pm$ 2.0s

### v2

@@ 6 Conclusion, limitations, and future work
We presented Trace, a new framework that combines contrastive learning with neighbor embeddings to directly generate interpretable 2D visualizations of large-scale neural time-series data. Using the inherent structure of multi-trial recordings common in neuroscience experiments, Trace is able to separate subtle differences of simulated neural response types that competing methods fail to distinguish. When applied to a diverse neural dataset of two-photon recordings, Trace captured both continuous variations in neural properties and discrete cell-type structures, and identified clusters with fine differences in functional responses and highlighted recording artifacts in the learned 2D representation (Sec. 5.3). Trace proved especially valuable for Neuropixels spike train data, because its approach of trial averaging preserves the underlying Poisson statistics better than standard augmentations used by other contrastive methods (e.g. temporal jitter or additive Gaussian noise).
@@
Conceptually, Trace has two advantages over existing methods. First, methods like $t$-SNE and UMAP based on neighborhood relationships in the high-dimensional space are sensitive to uninformative, noisy baseline activity and may not be able to detect short class-informative response periods (Sec. 5.1). Contrastive frameworks such as CEED rely on general-purpose data transformations and may mask subtle response differences between cell types. Second, Trace directly produces a 2D visualization while other contrastive methods require a separate reduction step. As a result, the embeddings produced by Trace were the most biologically interpretable as revealed by their ability to reflect key response characteristics (Sec. 5.2).
@@
A limitation of our work is the need to train an embedding model, increasing computational time compared to $t$-SNE or UMAP. However, this comes with the benefit of improved embedding quality and the ability to map new data points. Another limitation of our work is the need for repeated trials. However, this data collection approach data is standard for many neuronal time-series datasets. In some cases, between-trial variation may be of interest, e.g., when investigating adaptation over multiple trials. Trace operates under the assumption that differences between trials are undesirable noise thus making it not the right tool for this type of question.
In future work, it will be interesting to apply Trace to a much wider range of time-series data because many non-neuroscience datasets also have inherent multi-trial structure or can be reshaped into this format. For example, in sport analytics inertial measurements are often taken during repeated exercises, such as basketball free-throw drills [hoelzemann2023hang, garcia2022database]. The the Google Speech Command Dataset [warden2018speech] contains 5 audio recordings per speech commands and speaker. For medical data such as electrocardiography one could use different daily cycles or stress tests as trials. For financial market data, trading periods can be used as trials to detect unusual patterns or market anomalies. Another application domain could be climate data where years, seasons, or tidal cycles could be used as trials. These avenues highlight the broad applicability of leveraging multi-trial structure with Trace.
Acknowledgments
The authors thank the Hertie Foundation, the German Research Foundation (CRC 1233 “Robust Vision” and DFG/ANR EU 42/12-1, Project number 505379160) and the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for financal support. FS acknowledges the support of the Lower Saxony Ministry of Science and Culture (MWK) with funds from the Volkswagen Foundation’s zukunft.niedersachsen program (project name: CAIMed-Lower Saxony Center for Artificial Intelligence and Causal Methods in Medicine; grant number: ZN4257. KF was supported by the European Research Council (Starting Grant ”Eye To Action” 101117156). PB and DK are members of the EXC 2064 “Machine Learning — New Perspectives for Science“.
@@ Appendix A Experimental details
A.1 Identifying neural response groups for the calcium imaging dataset
A typical coarse classification of retinal ganglion cells is by their response to On- and Off-sets in the light stimulus and can typically be obtained with high confidence from simple statistics of the mean trial response of the moving bar stimulus [baden2016functional]. In particular, suppressed-by-contrast responses were identified as the ones with negative area under the curve, calculated from their mean response across all directions of the moving bar. ON-OFF responses were identified using the characteristic double-peak (2nd component in PCA captured this double-peak). Finally, ON and OFF responses were classified based on their response direction (captured in 1st component of PCA).
@@ A.2 ON-OFF index
Then, metric was clipped within the range [1,1][-1,1] to obtain a bounded metric.
@@ A.4 Outlier detection
A.5 Allen Institute Neuropixels spiking dataset post-processing
@@
A.6 Implementation details
Trace uses a lightweight multi-layer perceptron (MLP) that can run efficiently on a GPU. To ensure a fair comparison, we matched CEED’s architecture [vishnubhotla2023towards] and used an MLP consisting only of four layers with sizes [768, 512, 256, 128] and ReLU activations between them. We added a 1024-dimensional projection head. Due to TS2Vec’s specialized augmentations and hierarchical contrastive scheme, harmonizing its architecture was more difficult and we stuck to its original, dialated CNN architecture.
@@ A.6 Implementation details
For the calcium imaging dataset, we used optimal hyperparameters found with a grid search with batch sizes ranging from 1024 to 3200 and learning rates from 0.1 to 0.2 for both Trace and CEED. The best hyperparameter setting was chosen based on the final loss. For Trace the best batch size was 1280, while for CEED it was 1024. A learning rate of 0.1 was optimal for both. For Trace ++ CEED we used hyperparameters as in the unmodified Trace version.
For the Neuropixels dataset we ran a separate grid search and identified optimal hyperparameters for for Trace and CEED (0.03 learning rate, batch size 512) and for Trace ++ CEED (0.08 learning rate, batch size 1024).
[…]
Unless otherwise specified, we trained all embeddings for 1000 epochs for the calcium imaging dataset.
@@
A.7 CEED augmentations
Here we give details on the general-purpose time-series augmentaions that we used when running CEED and Trace++CEED. We apply these augmentation to the mean of all trials,
@@ A.7 CEED augmentations
$x^{\prime}_{i}=r^{\prime}\cdot x_{i}\,\text{ and }\,x^{\prime\prime}_{i}=r^{\prime\prime}\cdot x_{i}\quad\text{ with }r^{\prime},r^{\prime\prime}\sim\mathcal{U}(0.7,1.3),$ (10)
@@
$x^{\prime}_{i}[t]=x_{i}[t-s^{\prime}]\,\text{ and }\,x^{\prime\prime}_{i}[t]=x_{i}[t-s^{\prime\prime}]\quad\text{ for }s^{\prime},s^{\prime\prime}\sim\mathcal{U}\big(\{\pm 1,\pm 2,\pm 3\}\big).$ (11)
To efficiently handle the computationally expensive correlated noise generation, we pre-computed 10,000 noise samples from the covariance matrix of the data using a multivariate normal distribution and added them to the original time series, mimicking the augmentation
$x^{\prime}_{i}=x_{i}+\varepsilon^{\prime}\,\text{ and }\,x^{\prime\prime}_{i}=x_{i}+\varepsilon^{\prime\prime}\quad\text{ for }\varepsilon^{\prime},\varepsilon^{\prime\prime}\sim\mathcal{N}\big(\mathbf{0},\text{Cov}(\{x_{1},\dots,x_{n}\}\big).$ (12)
We applied each transformation independently, with probabilities of 0.7 for amplitude jittering, 0.6 for temporal jittering, and 0.5 for correlated noise injection. Hyperparameters of the transformations were adapted to our datasets.
A.8 Supplementary tables and figures
Figure S1: Trace is able to separate classes in the synthetic dataset even for high noise levels and outperforms other methods. (a) Trace embedding for increasing amounts of noise during the neural baseline activity. Discriminability (D) values indicated at the top of each embedding. Standard deviation (SD) of noise is indicated at the top of the figure. (b) Same as (a), but for CEED $\rightarrow$ PCA, (c) TS2Vec $\rightarrow$ PCA, (d) $t$-SNE, and (e) UMAP.
Table S1: Mean discriminability for simulated responses across three seeds for different noise levels (standard deviation of baseline response). Best performing method in bold.
@@ A.8 Supplementary tables and figures
Figure S2: Trace embeddings with varying numbers trials $k$ for averaging. (a) From left to right: Embeddings for the synthetic dataset with increasing numbers of trials $k$ from 1 to 7 used for the non-overlapping subsets. Discriminability (D) between class 1 and class 2 neurons was computed as in Fig. 2. Bottom: Loss over epochs. (b) From left to right: Embeddings for the calcium imaging dataset with increasing numbers of trials per subset mean $k$ from 1 to 7 for the ”chirp” stimulus and from 1 to 5 for the moving bars stimulus. Neurons are colored by the four broad groups OFF (blue), ON-OFF (turquoise), ON (green), Suppressed-by-contrast (black). Bottom: Loss over epochs.
Table S2: Reducing number of trials ($k$) used for the non-overlapping subsets using the calcium imaging dataset. The maximum number of trials was 7 for the ”chirp” stimulus ($k_{\text{chirp}}$) and 5 for the moving bar stimulus ($k_{\text{moving bar}}$). Best performance in bold.
@@
Figure S3: Trace-variants visually exhibit a good balance between resolving finer cluster structure and retaining large-scale structure of the calcium imaging dataset. (a – c) Two-dimensional embeddings of (a) Trace, (b) Trace+CEED, (c) Trace+TS2Vec. Color-coded according to their functional group.
@@
Figure S5: Identified artifacts in the calcium imaging dataset across all methods. (a) Trace embedding with identified artifact island in red. (b) Same as in (a) but for Trace ++ CEED, (c) Trace ++ TS2Vec, (d) CEED $\rightarrow$ PCA, (e) TS2Vec $\rightarrow$ PCA, (f $t$-SNE, and (g) UMAP.
@@
Table S3: Evaluating dimensionality reduction options for visualizing CEED’s 5-dimensional embeddings in 2D space. For the different normalizations, see the caption of Fig. S6. Best performance in bold. All metrics other than the aggregated rank are better when higher.
Model $r_{\text{OOi}}$ $r_{\text{RTi}}$ $r_{\text{Depth}}$ $\mu_{\text{Rank}}$
CEED (non-norm PCA) 0.58 0.14 0.54 2.33
CEED (non-norm UMAP) 0.36 0.19 0.54 3
CEED (unit-norm PCA) 0.55 0.11 0.55 3
CEED (unit-norm UMAP) 0.38 0.16 0.31 4.33
CEED (z-norm PCA) 0.59 0.12 0.53 3
CEED (z-norm UMAP) 0.42 0.09 0.39 5
@@
Table S4: Results for different methods on the calcium image data for varying embedding dimension. Trace performed best in terms of kkNN accuracy in all dimensions and maintained its performance in terms of Spearman correlation behind CEED, while TS2Vec degraded significantly. TS2Vec used the cosine similarity for $d>2$.
kkNN accuracy Spearman correlation $r_{S}$
Model $d=2$ $d=16$ $d=128$ $\mu_{\text{Rank}}$ $d=2$ $d=16$ $d=128$ $\mu_{\text{Rank}}$
Trace 69.7 75.3 75.5 1 0.45 0.4 0.41 2.33
CEED 64.0 72.3 72.2 2 0.51 0.58 0.51 1.33
TS2Vec 23.8 54.9 66.4 3 0.60 0.13 0.18 2.33
Table S5: Training time for all methods on the calcium imaging dataset. Data augmentations for CEED and TRACE ++ CEED were pre-computed and sampled from during training to improve training time (see Sec. A.7). Please note that no data augmentations are necessary for Trace. Each experiment was run on a NVIDIA A40 GPU 48 GB. For implementation details, see Sec. A.6.
@@
Trace ++ CEED 1000 25:40 $\pm$ 5.1s
Trace ++ TS2Vec 1000 962:43 $\pm$ 2668.8s
[…]